OpenStudy (anonymous):

Find the radius of convergence and interval of convergence of the following series.

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (-1)^n \frac{ (x-4)^n }{ 4n+1 }\]

OpenStudy (anonymous):

Set up the ratio test, but stuck from there \lim_{n \rightarrow \infty}|\frac{ (x-4)^n+1 }{ 4(n+1)+1 }*\frac{ 4n+1 }{ (x-4)^n }|\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}|\frac{ (x-4)^n+1 }{ 4(n+1)+1 }*\frac{ 4n+1 }{ (x-4)^n }|\]\]

OpenStudy (anonymous):

Found a similar example and got it - nevermind!